OpenStudy (anonymous):

Simplify: the square root of 36/121

OpenStudy (saifoo.khan):

\[\Large \sqrt{\frac{36}{121}} \to \frac{\sqrt{36}}{\sqrt{121}} \to \frac{6}{11}\]

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow \sqrt{36 \over 121} }\) \(\Large \color{purple}{\rightarrow {\sqrt{36} \over \sqrt{121}} }\)

Parth (parthkohli):

Solve now?

Parth (parthkohli):

Oh, saifoo solved it.

OpenStudy (anonymous):

Thanks!